A conical shaped tank of height H and width W is placed above a spherical water tank of radius R. Water drains from the conical tank into the spherical tank through a small opening of radius r. Similarly, the spherical tank empties though a small opening of radius r located at its bottom. Find the time it would take to drain both the tanks.
1
Expert's answer
2020-12-17T09:17:36-0500

Height of Cone = H

Width of cone = W

Width of cone = Circumference of cone = 2πR'

Base Radius of cone = W/2π

Volume of cone = 1/3πr²h

Volume of cone = "\frac{1}{3}×π×(\frac{W}{2π})^2×H = \frac{W²H}{12}"


Volume of spherical tank = 4/3πR³


Total volume of both objects = "\dfrac{W²H}{12π} + \dfrac{4πR³}{3} ="


"\dfrac{W²H + 16πR⁴}{12π}"


Since the radii of the openings of both the spherical and the conical tanks are equal, their rates are equal

The rate at which water moves away from the tanks = 2πr m³/sec


The time taken = volume/rate = "\dfrac{W²H + 16πR⁴}{12π} × \dfrac{1}{2πr} ="


"\dfrac{W²H + 16πR⁴}{24πr}"